OpenStudy (anonymous):

[\sqrt{d ^{2}-19} -2d + 11=0

OpenStudy (bahrom7893):

wait can u repost the question?

OpenStudy (bahrom7893):

is it the same that i answered?

OpenStudy (anonymous):

no i typed it wrong the first time

OpenStudy (bahrom7893):

So can u repost it, i still dont see the right question

OpenStudy (anonymous):

\[\sqrt{d ^{2}-19}-2d+11=0\]

OpenStudy (bahrom7893):

okay

OpenStudy (bahrom7893):

\[\sqrt{d^2 - 19} = 2d - 11\] Move -2d and 11 to the other side

OpenStudy (bahrom7893):

Square both sides: \[d^2 - 19 = (2d - 11)^2\]

OpenStudy (bahrom7893):

Rewrite the right side as: \[d^2 - 19 = 4d^2 - 44d + 121\]

OpenStudy (bahrom7893):

Rewrite in form of a regular quadratic equation and simplify: \[3d^2 - 44d +130 = 0\]

OpenStudy (bahrom7893):

Can u finish solving for d?

OpenStudy (anonymous):

yes, Do you have time for one more. It is similar with a cube root?

OpenStudy (bahrom7893):

yeah

OpenStudy (bahrom7893):

in that case u cube both sides, same method

OpenStudy (anonymous):

\[\sqrt[3]{8v ^{2}-6v}+1=0\]

OpenStudy (bahrom7893):

so move one over to the right and cube both sides to get: 8v^2 - 6v = (-1)^3 8v^2 - 6v = -1. Solve for V

OpenStudy (anonymous):

I get \[8v^{2}-6v+1\]

OpenStudy (bahrom7893):

yeah 8v^2 - 5v + 1 = 0

OpenStudy (bahrom7893):

solve for v

OpenStudy (anonymous):

thanks